# Constrained Error-Correcting Codes for Efficient DNA Synthesis

### v1

@@ IV Constrained ECCs for DNA synthesis
In this section, we construct the constrained ECCs for DNA synthesis, that not only satisfy the $\ell$-RLL and $\epsilon$-balance, but also are capable of correcting one indel in DNA sequences.
@@
respectively, where for each $i\in[1,n)$,
$s_{i}(\mathbf{c})=\left\{\begin{array}[]{lll}1&\text{~{}if~{}}c_{i+1}\geq c_{i},\\ 0&\text{~{}else.}\end{array}\right.$
Denote $m=n-\lceil\log_{4}n\rceil-3$. For any $\mathbf{c}=(c_{1},\dots,c_{m},c_{m+1},\dots,c_{n})\in\Sigma_{4}^{n}$, let $c_{m+4}\cdots c_{n}$ be the $4$-ary representation of $\mathrm{VT}(s(\mathbf{c}))\equiv a\bmod n$. Then, the systematic VT code constructed in [13] is
$\mathrm{VT}_{n}(a,b)=\left\{\mathbf{c}=(c_{1},\dots,c_{m},c_{m+1},\dots,c_{n})\in\Sigma_{4}^{n}:\right.$
$\mathrm{VT}(s(\mathbf{c}))\equiv a\bmod n,\qquad\qquad\quad\quad$
$c_{m+1}=c_{m+2}=c_{m}+2\bmod 4,\quad\quad$
$\left.c_{m+3}=\sum_{i=1}^{n}c_{i}\equiv b\bmod 4~{}\quad\quad\quad\right\}.$
Additionally, we define $\mathrm{VT}_{n}^{-1}(\mathbf{c},a,b)\triangleq(c_{1},c_{2},\dots,c_{m})\in\Sigma_{4}^{m}$.
For a symbol $a\in\Sigma_{4}$, let $f(a)=5-a$. Denote the length-$n$ constrained ECCs which satisfy $(\ell,\epsilon)$-constraints and can correct one indel in DNA synthesis by $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$. We propose the encoding and decoding methods of these constrained ECCs for DNA synthesis in algorithms 3 and 4, denoted by $\mathrm{Enc}_{T,\mathrm{indel}}$ and $\mathrm{Dec}_{T,\mathrm{indel}}$ respectively. In Algorithm 3, the original message $\mathbf{m}\in\{0,1\}^{k}$ is first encoded into a constrained codeword $\mathbf{c}^{\prime}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$, which is then concatenated with the parity symbols $\mathbf{p}$ of a systematic VT code to form the codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$. To ensure the $(\ell,\epsilon)$-constraints, additional symbols $\beta$ and $f(\cdot)$ are introduced. In Algorithm 4, the received vector $\mathbf{c}$ is first decoded by determining the length of $\mathbf{c}$ and then applying the VT decoder. Finally, the original message is recovered by decoding the constrained code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$.
Algorithm 3 $\mathrm{Enc}_{T,\mathrm{indel}}$: Encoder of constrained ECCs for DNA synthesis
@@
0:  $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$.
@@
Algorithm 4 $\mathrm{Dec}_{T,\mathrm{indel}}$: Decoder of constrained ECCs for DNA synthesis
@@
1:  If $|\mathbf{c}|=n$, no indel happened, then $\mathbf{c}^{\prime}=(c_{1},\dots,c_{k_{1}})$, where $k_{1}=k+r_{\ell,\epsilon,T_{1}}$. Return $\mathbf{m}=\mathrm{Dec}_{\ell,\epsilon,T_{1}}(\mathbf{c}^{\prime})$.
@@
For given $n,\ell,\epsilon$ and a fixed $n\leq T\leq 4n$, denote $T_{1}=T-4(2\lceil\log_{4}n\rceil+8)$. The code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$ encodes a binary sequence with length $\lceil\log_{2}{N(n,\ell,\epsilon,T_{1})}\rceil$ to a codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$, which can correct one indel in DNA sequences. This results in a redundancy of $r_{\ell,\epsilon,T_{1}}+2\lceil\log_{4}n\rceil+8$. Additionally, the encoding and decoding algorithms run in $O(n^{4}\ell^{2})$ time.
@@
The properties of indel-correcting, $(\ell,\epsilon)$-constraints, along with the number of redundancy of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$ have been clearly established. We now focus on the computation complexity and synthesis time. Since the encoder and decoder of systematic VT codes are linear, then the computation complexity of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$ is the same as that of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$, i.e., $O(n^{4}\ell^{2})$.
In addition, the synthesis time of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$ is given by $T_{1}=T-4(2\lceil\log_{4}n\rceil+8)$. Therefore, the synthesis time of a codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,B^{DI})$ is an most $T$, since the additional redundancy is $82\lceil\log_{4}n\rceil+$ 8 symbols. This completes the proof. ∎
@@ V Conclusions
In this paper, we propose capacity-achieving constrained codes with either only $\ell$-RLL or both $\ell$-RLL and $\epsilon$-balanced constraints for DNA synthesis. Furthermore, we extend these results to design constrained ECCs that can correct a single indel for DNA synthesis. Note that it is also possible to construct the constrained codes with only $\epsilon$-balance, or the constrained ECCs capable of correcting an edit error using a similar approach.

### v3

@@ IV Constrained ECCs for DNA Synthesis
In this section, we construct the constrained ECCs for DNA synthesis, that not only satisfy the $\ell$-RLL and $\epsilon$-balance constraints, but also are capable of correcting one indel in DNA sequences.
@@
respectively, where for each $i\in[1,n)$, $s_{i}(\mathbf{c})=1$ if $c_{i+1}\geq c_{i}$, otherwise, it is 00.
Denote $m=n-\lceil\log_{4}n\rceil-3$. For any $\mathbf{c}=(c_{1},\dots,c_{m},$ $c_{m+1},\dots,c_{n})\in\Sigma_{4}^{n}$, let $c_{m+4}\cdots c_{n}$ be the $4$-ary representation of $\mathrm{VT}(s(\mathbf{c}))\equiv a\bmod n$. Then, the systematic VT code constructed in [13] for $4$-ary sequences is
$\mathrm{VT}_{n}(a,b)=\Big{\{}\mathbf{c}=(c_{1},\dots,c_{m},c_{m+1},\dots,c_{n})\in\Sigma_{4}^{n}:$
$\mathrm{VT}(s(\mathbf{c}))\equiv a\bmod n,$
$c_{m+1}=c_{m+2}=c_{m}+2\bmod 4,$
$c_{m+3}=\sum_{i=1}^{m}c_{i}\equiv b\bmod 4\Big{\}}.$
Moreover, we define $\mathrm{VT}_{n}^{-1}(\mathbf{c},a,b)\triangleq(c_{1},c_{2},\dots,c_{m})\in\Sigma_{4}^{m}$.
For a symbol $a\in\Sigma_{4}$, let $f(a)=5-a$. Denote the length-$n$ constrained ECC which satisfies $(\ell,\epsilon)$-constraints and can correct one indel in DNA synthesis by $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$. We propose the encoding and decoding methods of the constrained ECC for DNA synthesis in Algorithms 3 and 4, denoted by $\mathrm{Enc}_{T,\mathrm{indel}}$ and $\mathrm{Dec}_{T,\mathrm{indel}}$, respectively. In Algorithm 3, the original message $\mathbf{m}\in\{0,1\}^{k}$ is first encoded into a constrained codeword $\mathbf{c}^{\prime}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$, which is then concatenated with the parity symbol $\mathbf{p}$ of a systematic VT code to form the codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$. To ensure the $(\ell,\epsilon)$-constraints, additional symbols $\beta$ and $f(\cdot)$ are introduced. In Algorithm 4, the received vector $\mathbf{c}$ is first decoded by determining the length of $\mathbf{c}$ and then applying the VT decoder. Finally, the original message is recovered by decoding the constrained code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$.
Algorithm 3 $\mathrm{Enc}_{T,\mathrm{indel}}$: Encoder of the designed constrained ECC for DNA synthesis.
@@
0:  $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$.
@@
Algorithm 4 $\mathrm{Dec}_{T,\mathrm{indel}}$: Decoder of the designed constrained ECC for DNA synthesis.
@@
1:  If $|\mathbf{c}|=n$, no indel occurred, then $\mathbf{c}^{\prime}=(c_{1},\dots,c_{k_{1}})$, where $k_{1}=k+r_{\ell,\epsilon,T_{1}}$. Return $\mathbf{m}=\mathrm{Dec}_{\ell,\epsilon,T_{1}}(\mathbf{c}^{\prime})$.
@@
For given $n,\ell,\epsilon$ and a fixed $T$ with $n\leq T\leq 4n$, denote $T_{1}=T-4(2\lceil\log_{4}n\rceil+8)$. The code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$ encodes a binary sequence with length $\lceil\log_{2}{N(n,\ell,\epsilon,T_{1})}\rceil$ to a codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$, which can correct one indel in DNA sequences. This results in a redundancy of $r_{\ell,\epsilon,T_{1}}+2\lceil\log_{4}n\rceil+8$. Additionally, the encoding and decoding algorithms run in $O(n^{4}\ell^{2})$ time.
@@
The properties of $(\ell,\epsilon)$-constraints, indel-correcting, along with the number of redundancy of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$ have been clearly established. We now focus on the computational complexity and synthesis time. Since the encoder and decoder of the systematic VT codes constructed in [13] for $4$-ary sequences have linear time complexity, then the computational complexity of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$ is the same as that of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$, i.e., $O(n^{4}\ell^{2})$.
In addition, the synthesis time of $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T_{1})$ is given by $T_{1}=T-4(2\lceil\log_{4}n\rceil+8)$. Therefore, the synthesis time of a codeword $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T,\mathcal{B}^{DI})$ is at most $T$, since the additional redundancy is $82\lceil\log_{4}n\rceil+$ 8 symbols. This completes the proof. ∎
@@ V Conclusions
In this paper, we propose capacity-achieving constrained codes with either only $\ell$-RLL or both $\ell$-RLL and $\epsilon$-balanced constraints for DNA synthesis. Furthermore, we extend these results to design constrained ECCs that can correct a single indel for DNA synthesis. Note that it is also possible to construct constrained codes with only $\epsilon$-balance, or ECCs capable of correcting an edit error using a similar approach.